Firm A produces bags and expects to introduce a new product in a market. Firm A adopts indirect distribution channel in the order of Firm A → Wholesaler → Retailer → Buyer (or Consumer).

Q1. If a buyer purchases the new product for $109.99 at retailer, what is unit price did firm A sell the new product to its wholesaler if a 10% mark up on sales on every new product is applied to each channel member? Let's assume that wholesaler's final unit cost is 5% higher than its acquisition unit cost and the retailer's final unit cost is 10% higher than its acquisition unit cost.

Q2. If the break-even volume is 40,000 units for Firm A, what is Variable Cost per unit at break-even. Let's assume that Firm A's Total Fixed Costs are $2,000,000.

Q3. If Firm A anticipates that 300,000 new products will be sold through channel members' sales, how much is total profit? (Tip: Think that FIrm A's Total Fixed Costs are $2,000,000, break-even is 40,000 units, so 300,000 units at break-even is not effective anymore)

Q4. What is the price per unit that Firam A can't sell the new products profitably to its wholesaler? Although the wholesaler accepts Firm A's competitor's products, the wholesaler still has the committment to sell 300,000 units of Firm A's product.

Q5. What can be an optimal pricing strategy for Firm A if it is in a matured market?
Answer #1
  1. Given retailer to buyer the price is 109.99, the price for retailer from the wholesaler is=109.99/(1+10%)=$99.99
    Firm A to wholesaler it is =99.99/(1+10%)=$90.90
    Answer is $90.90
  2. Let variable costs be “x” per unit then
    Fixed costs= sales-variable costs( breakeven state)
    2000000=40000*(90.9-x)
    X=$40.90
  3. Profit= sales-variable-fixed
    =(300000*(90.9-40.90))-200000=$14,800,000